Pacific Ink had beginning work-in-process inventory of $760,000 on October 1. Of this amount, $305,920 was the cost of direct materials and $454,080 was the cost of conversion.The 49,000 units in the beginning inventory were 30 percent complete with respect to both direct materials and conversion costs. During October, 104,000 units were transferred out and 31,000 remained in ending inventory.The units in ending inventory were 80 percent complete with respect to direct materials and 40 percent complete with respect to conversion costs. Costs incurred during the period amounted to $2,282,000 for direct materials and $3,101,850 for conversion. 3. value: 14.28 points Required information

a-1. Compute the cost of goods transferred out and the cost of ending inventory using the FIFO method. (Round intermediate calculations to 2 decimal places.)
Answer #1

Solution

Total
Units Completed $ 5,269,650
Ending Inventory $ 874,200

Working

STEP 1 (FIFO)
Reconciliation of Units
Units % already completed % completed this period
Material Conversion cost Material Conversion cost
Beginning WIP             49,000 30% 30% 70% 70%
Units introduced             86,000
Total units to be accounted for           135,000
Completed and Transferred unit           104,000 0% 0% 100% 100%
Ending WIP             31,000 0% 0% 80% 40%

.

STEP 2 (FIFO)
Equivalent Units
Total Units Material Conversion cost
Units Transferred: % completed this period Equivalent Units % completed this period Equivalent Units
From WIP             49,000 70%                   34,300 70%                        34,300
From units started/Introduced             55,000 100%                   55,000 100%                        55,000
Total           104,000                   89,300                        89,300
Total Units Material Conversion cost
Ending WIP % completed this period Equivalent Units % completed this period Equivalent Units
Total             31,000 80%                   24,800 40%                        12,400
Total EUP           135,000                 114,100                     101,700

.

STEP 3
Cost per Equivalent Units AND Cost to be accounted for
Material Conversion cost TOTAL
Cost incurred in Current Period $ 2,282,000 $ 3,101,850 $ 5,383,850
Total Equivalent Units                                            114,100                                                     101,700
Cost per Equivalent Units $ 20.00 $ 30.50 $ 50.50
Cost of Beginning WIP $ 760,000
Total cost to be accounted for $ 6,143,850

.

STEP 4
Cost report [Including Reconciliation]
Cost of Units Transferred            104,000.00 units
Equivalent Units [Step 2] Cost per Equivalent Units [Step 3] Cost of Units Transferred TOTAL
A From Beginning WIP:
Material                  34,300 $ 20.00 $ 686,000
Conversion cost                  34,300 $ 30.50 $ 1,046,150 $ 1,732,150
B From units started & completed
Material                  55,000 $ 20.00 $ 1,100,000
Conversion cost                  55,000 $ 30.50 $ 1,677,500 $ 2,777,500
C=A+B Total $ 4,509,650
D Cost of Beginning WIP $ 760,000
E=C+D Total Cost of Units transferred $ 5,269,650
Cost of Ending WIP               31,000.00 units
Equivalent Units [Step 2] Cost per Equivalent Units [Step 3] Cost of Ending WIP TOTAL
F Material                  24,800 $ 20.00 $ 496,000
G Conversion cost                  12,400 $ 30.50 $ 378,200
H = F + G Total Cost of Ending WIP $ 874200.0
I = E + H Total Cost accounted for $ 6,143,850
J Rounding Off Differences $ 0
K = I + J Total cost to be accounted for [equals to computed in STEP 3] $ 6,143,850